OpenStudy (adunb8):

anyone know how to factor x^2+x+4?

OpenStudy (anonymous):

if im not wrong its prime!

OpenStudy (adunb8):

? why are you asking about \[x^4?\]

OpenStudy (anonymous):

Sorry, misread your question. It cannot be factor.

OpenStudy (adunb8):

how so?

OpenStudy (adunb8):

oh sorry this is actually linear algebra question using cauchy euler equation

OpenStudy (anonymous):

Use the quadratic formula by plugging in the variables from the given equation. So taking your given equation.. \[x ^{2}+x+4\] a=1 b=1 c=4 according to... \[ax ^{2}+bx+c = 0\] Plug your numbers accordingly and you will get your answer when you set to zero. \[\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\] You should get... \[\frac{ -1+\sqrt{15} }{ 2 },\frac{ -1-\sqrt{15} }{ 2 }\]

OpenStudy (anonymous):

You can't factor this equation.

OpenStudy (sirm3d):

the trinomial cannot be factored. expression that cannot be factored is called prime.